OpenStudy (anonymous):

I AM SO LOST: Solve with matrices. 5x -3y = 18 and 2x + 7y = -1 Indicate each matrix with the vertical bar (|) symbol from the keyboard at the beginning and end of each row.

OpenStudy (anonymous):

\[\left[\begin{matrix}5 & -3 \\ 2 & 7\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}18 \\ -1\end{matrix}\right)\] or AX=B 

OpenStudy (anonymous):

your my hero got it thanks